OpenStudy (anonymous):

Find the solution of y = 6x + 1 for x = 5.

OpenStudy (merchandize):

put the value of x=5 in the equation.. y=6x+1 y=6(5)+1 y=30+1 y=31

OpenStudy (anonymous):

y= 6(5) + 1 y= 30 + 1 y= 31 Solution Set {5, 31}

OpenStudy (anonymous):

thnk you guys

OpenStudy (anonymous):

no problem... :)